OpenStudy (anonymous):

What is the difference between 'x' being proportional to 'y' and 'y' being proportional to 'x'??

OpenStudy (anonymous):

(X-y)(x+y)/xy

OpenStudy (unklerhaukus):

The proportionality constants will be inverses: \[x\propto y\implies x=\alpha y\\[2ex] y\propto x\implies y=\beta x\\[2ex] \therefore\quad \alpha =\frac 1\beta \]